A population of flour beetles has 1000 individuals. Normally the beetles are red in color; however, this population has a utation in the gene for body color that results in black beetles that are designated by the rr genotype. Red body color is dominant to black so the genotypes RR and Rr result in red beetles. Assume the population is in Hardy-Weinberg equilibrium, with a 50% frequency of each allele for body color.

A) --- what is the frequency of the recessive genotype, rr? _______

   --- the homozygous dominant genotype RR? _________

     --- the heterozygous genotype Rr? ___________

B) --- How many moths in the population are black? __________

     --- How many moths in the population are red? _________

C) --- What would be the percentages of the body color phenotypes after 50 generations if the population  

          maintains the conditions of Hardy Weinberg equilibrium?

     --- percentage of black phenotype?

     --- percentage of red phenotype?

D) what would the allele frequencies be if a populaiton bottleneck occurred and only 5 individuals survived, two female re heterozygotes and 3 black males?

---Frequency of R?_______

---Frequency of r?________

E) If the population in (d) randomly mated, what would be the allele frequencies of their offspring?

--- frequency of R?_____

---frequency of r?_______
Answer #1

A) Frequency of r allele is 0.5

Frequency of R allele is 1-r =1- 0.5=0.5

Frequency of rr is 0.5×0.5=0.25

Frequency of homozygous dominant RR is 0.5×0.5=0.25

Frequency of heterozygous Rr is 2×R×r = 2×0.5×0.5=0.50

B)  

Number of black moth are 0.25×1000=250

Number of Red moth are 0.75×1000= 750

C)

The genotype frequency remains constant from generation to generation.

Percentage of Black = 0.25×100= 25%

Percentage of red are = 100-25=75%

D)

Frequency of r is = genotype frequency of ( black + (1/2) × heterozygous)

= (3/5) +(1/2) × 2/5

= 0.6+ 0.2= 0.8

Allelic frequency of R is 1-r= 1-0.8=0.2

E) The allelic frequency remains constant in offsprings. That's why frequency of R is 0.2

And allelic frequency of r is 0.8